OpenStudy (shubhamsrg):

x,y,z are non zero real numbers. a,b,c are complex numbers such that |a| = |b| = |c| =1 . If x+y+z = 0 = ax + by + cz , how do I prove a=b=c ?

OpenStudy (experimentx):

\[ (x+y)^2 = z^2 \] complex numbers add up like vectors, write \( a = \cos \theta_1 + i \sin \theta_1, \) ... and so on. \[ ax+by+cz = 0 \\ ax+by = -cz \\ (x \cos \theta_1 + y \cos \theta_2 ) + i (x \sin \theta_1 + y \sin \theta_2 ) = z( \cos \theta_3 + i \sin \theta_3) \\ |(x \cos \theta_1 + y \cos \theta_2 ) + i (x \sin \theta_1 + y \sin \theta_2 )| = |z( \cos \theta_3 + i \sin \theta_3)| \\ x^2 + 2xy ( \cos \theta_1 \cos \theta_2 +\sin \theta_1 \sin \theta_2) + y^2 = z^2 \\ x^2 + 2 xy \cos(\theta_1 - \theta_2) +y^2 = z^2 = (x+y)^2\]

OpenStudy (experimentx):

that implies \(\theta_1 = \theta_2\) next time we will switch variables and prove \(\theta_1 = \theta_3\) ... hence the argument of a,b,c are equal

OpenStudy (shubhamsrg):

yes that was easy and i should have been able to do that hmm thanks again sire.

OpenStudy (experimentx):

yw